# Testing Dark Matter with Generative Models for Extragalactic Stellar Streams

### v1

@@ Abstract
1 Introduction
The nature of the dark matter particle, e.g., its free-streaming velocity, its mass, and whether it self-interacts, affects the overall properties of dark matter halos. Under $\Lambda$CDM, the density profiles of dark matter halos is expected to follow a Navarro–Frenk–White profile (NFW; Navarro et al. 1996), with a steeply rising cusp at the center and a characteristic falloff in the outskirts. Self-interacting dark matter (SIDM) models predict that the inner density profile of halos is cored with a constant density, while the outer part follows an NFW profile (Spergel & Steinhardt, 2000). Warm dark matter models can produce cores in their centers, but some fine-tuning is required (e.g., Strigari et al., 2007; Macciò et al., 2012) to match the observed cores in some galaxies. Fuzzy dark matter models predict a central soliton core, resulting in a flat inner density slope, while the outer halo transitions to an NFW-like profile (Hui et al., 2017). Therefore, mapping the variations in density slopes can provide insights into the nature of dark matter.
Observations of stellar kinematics and HI in gas rich dwarfs and spiral galaxies show examples of near linearly rising rotation curves, which implies that they have cored density profiles in their centers (e.g., Flores & Primack, 1994; Moore, 1994; Gentile et al., 2004; Oman et al., 2015; Oh et al., 2015; Lelli et al., 2016; Read et al., 2017; Hayashi et al., 2025). The fact that some dwarf spheroidal galaxies host globular clusters at present day can also be used to argue in favor of cored density profiles, as cored profiles would allow for globular clusters to remain close to their currently observed locations for long times instead of in-spiraling due to dynamical friction (e.g. in Fornax: Goerdt et al., 2006; Cole et al., 2012). $\Lambda$CDM predicts steeply rising cusps at the center of galaxies, thus these observations pose a challenge to the theory.
Williams et al. (2025) recently suggested a theoretical model based on statistical mechanics which shows that dark matter halos can have a diverse range of density profiles, including cored density cores at their centers. The discrepancy between their model and the lack of cores in dark matter only $\Lambda$CDM $N$-body simulations remains a mystery. Including baryons to DM simulations shows that galaxies could have more cored profiles even in $\Lambda$CDM due to bursty star formation and stellar feedback (see e.g., Oman et al., 2015; Straight et al., 2025), but the diversity of radial profiles of dark matter halos remain challenging to reproduce in simulations. Inferring cores from rotation curves involves assumptions about gas pressure (Pineda et al., 2017), and while alternative methods of inferring density slopes without assumptions about dynamical equilibrium exist (e.g., Nguyen et al., 2023), the sample sizes of measured cored density profiles remain low.
@@ 1 Introduction
The outer halo density profile (i.e., beyond the halo scale radius) depends on the overall amount of accretion experienced by a halo over the past few billion years (Diemer & Kravtsov, 2014). Outer halo slopes are also shaped by filamentary accretion from the cosmic web (e.g., Zhang et al., 2013; Euclid Collaboration: Laigle et al., 2025; Arora et al., 2025) and show a correlation with varying environment, where isolated halos have shallower outer density slopes (Avila-Reese et al., 1999). Although steepening of outer halo profiles can be detected with weak-lensing analyses of massive clusters (Diemer & Kravtsov, 2014), measuring the outer radial profiles of individual halos remains a challenge.
@@
Stellar streams offer an independent gravitational tracer of dark matter halos. When stellar streams form, stars escape from their progenitors due to the underlying mass profile of the host galaxy. The subsequent trajectories of stars around their host galaxy is typically dominated by the dark matter halo properties of the host Johnston et al. (2001). Stellar streams thus encode properties of dark matter halos. Studies of stellar streams in the Milky Way have shown that streams are sensitive to the local acceleration field (Bonaca & Hogg, 2018; Nibauer et al., 2022; Nibauer & Bonaca, 2025), and the shape of the dark matter halo (e.g., Koposov et al., 2010; Law & Majewski, 2010; Pearson et al., 2015; Küpper et al., 2015; Bovy et al., 2016; Yavetz et al., 2021, 2023; Woudenberg & Helmi, 2024). We are now entering an era where it will be possible to extend studies of the Milky Way’s dark matter halo to thousands of other galaxies with upcoming low surface brightness observations from Euclid (Racca et al., 2016), The Rubin Observatory (Ivezić et al., 2019), The Nancy Grace Roman Space Telescope (Roman; Spergel et al. 2015), and ARRAKIHS (Guzman et al., 2022).
@@
One of the first detections of an extragalactic stellar stream was the stream north of M83 shown in Malin & Hadley (1997) and discussed in van den Bergh (1980). We now know of more than 150 extragalactic streams from imaging individual galactic stellar halos (e.g., Mart´ınez-Delgado et al., 2008; McConnachie et al., 2009; Crnojević et al., 2016), and from larger surveys such as DESI (Mart´ınez-Delgado et al., 2023), DES (Miró-Carretero et al., 2024a), SAGA (Miró-Carretero et al., 2024a), HSC (Kado-Fong et al., 2018), and MADCASH (Carlin et al., 2016, 2019)). Already we are in a regime where the number and diversity of tidal features is applicable to dynamical modeling.
@@
Pearson et al. (2022b) developed an extragalactic stream-fitting code that explored a grid of 2D velocity vectors for the progenitor across ten halo mass bins using particle spray simulations (Fardal et al., 2015). They evaluated model accuracy by comparing stream tracks to observed control points and ensured realistic tidal stripping using a Jacobi radius criterion. Results were validated with follow-up $N$-body simulations (Kawata & Gibson, 2003). Applying their method to a stream near Centaurus A (Crnojević et al., 2016), they showed that a single progenitor radial velocity can constrain the dark matter halo mass ($M_{200}>4.7\times 10^{12}~M_{\odot}$). Additionally, radial velocity and distance gradients along the stream offer further mass constraints. However, a more exhaustive parameter search over both progenitor and halo properties is still needed, and was not performed in Pearson et al. (2022b) due to computational limitations of their approach.
@@
In spherical dark matter halo potentials, where the orbital plane itself does not precess, the precession per orbit between apocenter or pericenter is determined by both the orbital properties (e.g., ratio of radial to azimuthal periods; Johnston et al. 2001; Binney & Tremaine 2008; Hendel & Johnston 2015), and the slope of the potential (Belokurov et al., 2014). Walder et al. (2024) recently showed that for an orbit launched at apocenter and viewed at a 0 degree inclination, multiple wraps of the orbit can constrain the radial profile of the dark matter halo.
@@
The paper is organized as follows: In §2, we provide an overview of stellar stream evolution in potentials with different radial profiles. In §3, we introduce our method and lay out the details of the sampling technique. In §4, we show the results of applying our method to distinct streams on different orbits, in §5 we discuss our findings, and in §6 we summarize and conclude.
2 Stream morphologies as a probe of radial profiles
@@
where $M_{\rm halo}$ is the halo scale mass, $r_{s}$ is the scale radius of the potential, $c_{\rm NFW}$ is the halo concentration, $\gamma$ is the slope of the inner density profile, and $\beta$ is the slope of the density profile beyond the scale radius. The exponential cutoff ensures a finite mass. We use $r_{\rm cut}=25\times r_{s}$, and solve Poisson’s equation on a radial grid to infer an interpolated potential from the density. This choice of $r_{\rm cut}$ is not crucial to our modeling, since all streams have apocenters much lower than the minimum $r_{\rm cut}$ used (i.e., for $r_{s,\rm{min}}=10~\rm{kpc}$, $r_{\rm cut}=250~\rm{kpc}$). In Fig. 1 we show examples of three different dark matter halo density profiles: (1) An NFW profile (Navarro et al., 1997) with $(\gamma$,$\beta)$ = (1,31,31 , 3; solid line), (2) a cored profile, with $(\gamma$,$\beta)$ = (0.2,30.2,30.2 , 3; dotted line), and (3) an inner slope following an NFW profile, but a shallower outer slope $(\gamma$,$\beta)$ = (1,21,21 , 2; dashed line).
@@
Table 1: Parameters for the inner and outer stream
@@
footnotetext: Note that $r_{\rm peri}$ and $r_{\rm apo}$ are listed for the six different streams shown in Fig. 2 evolved in the three different potentials shown in Fig. 1. The first values for each stream are for the fiducial NFW halo.
@@
To illustrate the sensitivity of streams to variations in halo density profiles, in Fig. 2 we show two different streams which we label the inner stream (top row) and the outer stream (bottom row). The streams are evolved in the three different potentials shown in Fig. 1 (left: NFW, middle: cored, right: shallower outer profile), with a halo scale mass defined in Eq. 2 of $M_{\rm halo}=10^{13}$ M (corresponding to $M_{\rm enc}$ at the outer stream’s radius of $\approx 4\times 10^{12}$), $r_{\rm s,halo}=22$ kpc, $c_{\rm NFW}=15$, including a disk (Miyamoto & Nagai, 1975) with $m_{\rm disk}=5\times 10^{10}$ M, $a_{\rm disk}=3$ kpc, $b_{\rm disk}=0.2$ kpc. Our halo mass is higher than typical spiral galaxies in the local group, though is reasonable given lower limits on $M_{200}$ from Centaurus A (Pearson et al., 2022b) [^1] Note that the density structure of faint dwarf galaxies ($M_{*}\sim 10^{6}$ M) would be of particular interest, as stellar feedback in such galaxies would not modify the central structure (Bullock & Boylan-Kolchin, 2017).. We list the initial conditions and properties of each of the streams in Table 1. Note that throughout the paper, we fix the disk parameters, but we vary the halo parameters (see Table 2). The red stars in the upper panels show the progenitor location, and the red dashed line shows the past 500 Myr of evolution for each progenitor. Note that the streams were evolved for 5 Gyr.
@@
$\begin{split}\hat{v}_{x}&=\sin(\theta)\cos(\phi),\\ \hat{v}_{y}&=\sin(\theta)\sin(\phi),\\ \hat{v}_{z}&=\cos(\theta),\\ {\bf v}&=v\hat{v}.\end{split}$ (3)
@@
The top two panels of Fig. 3 show the angle between successive apocenters (i.e., the angle between orbital turning points), $\theta_{\rm apo}$, for the streams evolved under the mass model in Eq. 1 with varying $\gamma$ and $\beta$ (see labels) as a function of galactocentric radius. We calculate $\theta_{\rm apo}$ using the epicycle approximation (Appendix A). In Fig. 3, steeper inner and outer density slopes (larger $\gamma$ and larger $\beta$) lead to smaller angles between successive apocenters. Note, however, that at small galactocentric radii $\theta_{\rm apo}$ is more sensitive to changes in $\gamma$, whereas at larger galactocentric radii $\theta_{\rm apo}$ is more sensitive to changes in $\beta$ due to the relations between the radial and azimuthal periods (see Appendix A). We can understand the trends between $(\gamma,\beta)$ and $\theta_{\rm apo}$ as follows. As $\gamma$ increases, the inner regions of the potential become deeper and the radial and azimuthal periods both decrease, with their ratio approaching unity. Thus, the angle between apocenters diminishes. For $\beta$, the mass enclosed approximately goes like $M_{\rm enc}\propto r^{3-\beta}$ below the cutoff radius, $r<r_{\rm cut}$, though beyond the scale radius, $r_{s}$. At large $r$, increasing to $\beta>3$ the mass distribution becomes more centrally packed-in and orbits become approximately Keplerian, for which $\theta_{\rm apo}\xrightarrow[]{}0$. Thus, we expect that increasing $\beta$ causes turning point angles to decrease. Indeed, this is what we find in Fig. 2, and show analytically in Fig. 3. We note that the angles between turning points are degenerate with projection effects. The generative method we develop in §3 will account for viewing angle degeneracies.
@@
We have demonstrated that two different extragalactic streams at two different locations in their host halo are sensitive to different parts of the potential (as expected from Milky Way studies, e.g., Bonaca & Hogg, 2018; Nibauer et al., 2022), and that streams, which cross or are within the scale radius, are particularly sensitive to the inner slope of the potential. Note that while deviations from spherical symmetry can also introduce orbital plane precession, in this paper we focus on spherical dark matter halos as the simplest mass model that is also interesting for dark matter physics (Spergel & Steinhardt, 2000; Bullock & Boylan-Kolchin, 2017).
@@ 3 Method
3.1 Stream Generation and Model Parameters
To generate stellar streams in different potential models, we use the particle-spray method (Fardal et al., 2015) implemented in the GPU accelerated code streamsculptor (Nibauer et al., 2025), built within the JAX python framework (Bradbury et al., 2018). The GPU support of our simulator allows us to efficiently generate stream models in batches of 500500500, where each stream in the batch is simulated within a different underlying potential. Each model stream consists of 1000 particles, which is sufficient for the low resolution data representation that we will discuss in §3.2. Despite the lack of dynamical friction and star-by-star force calculations, the particle spray approach successfully characterizes the tracks of dwarf galaxy streams when compared to $N-$body simulations (Fardal et al., 2015; Pearson et al., 2022b).
We consider the two streams presented in §2, which trace different locations of the same potential. We aim to investigate whether the full radial profile of the host halo (i.e., the inner and outer density slopes) can be recovered from the morphology of the two streams. In Table 2 we list all the free and fixed parameters adopted in our sampling.
To generate streams, we fix the on-sky location of the progenitors. Several extragalactic streams have tentative progenitor locations (see e.g., Pearson et al. 2022b; Miró-Carretero et al. 2024b). However, future work can add free parameters for the progenitor’s on-sky location (e.g., along the ridgeline of the stream’s track). We find that incorrectly choosing the angular location of the progenitor biases progenitor parameters, but does not change our conclusions on inferring the full radial density profile (inner and outer) of the halo. We fit for the line-of-slight location of the progenitor in its host galaxy, $y_{\rm prog}$, along with its present-day velocity direction, sampled on the velocity unit sphere (see §2 and Table 2). When sampling from our prior, we test whether the progenitor’s orbit is bound. We reject parameters that produce unbound orbits.

### v2

@@ Abstract
I Introduction
[…]
@@ I Introduction
The outer halo density profile (i.e., beyond the halo scale radius) depends on the overall amount of accretion experienced by a halo over the past few billion years (Diemer and Kravtsov, 2014). Outer halo slopes are also shaped by filamentary accretion from the cosmic web (e.g., Zhang et al., 2013; Euclid Collaboration: Laigle et al., 2025; Arora et al., 2025) and show a correlation with varying environment, where isolated halos have shallower outer density slopes (Avila-Reese et al., 1999). Although steepening of outer halo profiles can be detected with weak-lensing analyses of massive clusters (Diemer and Kravtsov, 2014), measuring the outer radial profiles of individual halos remains a challenge.
@@
Stellar streams offer an independent gravitational tracer of dark matter halos. When stellar streams form, stars escape from their progenitors due to the underlying mass profile of the host galaxy. The subsequent trajectories of stars around their host galaxy is typically dominated by the dark matter halo properties of the host Johnston et al. (2001). Stellar streams thus encode properties of dark matter halos. Studies of stellar streams in the Milky Way have shown that streams are sensitive to the local acceleration field (Bonaca and Hogg, 2018; Nibauer et al., 2022; Nibauer and Bonaca, 2025), and the shape of the dark matter halo (e.g., Koposov et al., 2010; Law and Majewski, 2010; Pearson et al., 2015; Küpper et al., 2015; Bovy et al., 2016; Yavetz et al., 2021, 2023; Woudenberg and Helmi, 2024). We are now entering an era where it will be possible to extend studies of the Milky Way’s dark matter halo to thousands of other galaxies with upcoming low surface brightness observations from Euclid (Racca et al., 2016), The Rubin Observatory (Ivezić et al., 2019), The Nancy Grace Roman Space Telescope (Roman; Spergel et al. 2015), and ARRAKIHS (Guzman et al., 2022).
@@
One of the first detections of an extragalactic stellar stream was the stream north of M83 shown in Malin and Hadley (1997) and discussed in van den Bergh (1980). We now know of more than 150 extragalactic streams from imaging individual galactic stellar halos (e.g., Martínez-Delgado et al., 2008; McConnachie et al., 2009; Crnojević et al., 2016), and from larger surveys such as DESI (Martínez-Delgado et al., 2023), DES (Miró-Carretero et al., 2024b), SAGA (Miró-Carretero et al., 2024b), HSC (Kado-Fong et al., 2018), and MADCASH (Carlin et al., 2016, 2019)). Already we are in a regime where the number and diversity of tidal features is applicable to dynamical modeling.
@@
Pearson et al. (2022b) developed an extragalactic stream-fitting code that explored a grid of 2D velocity vectors for the progenitor across ten halo mass bins using particle spray simulations (Fardal et al., 2015). They evaluated model accuracy by comparing stream tracks to observed control points and ensured realistic tidal stripping using a Jacobi radius criterion. Results were validated with follow-up $N$-body simulations (Kawata and Gibson, 2003). Applying their method to a stream near Centaurus A (Crnojević et al., 2016), they showed that a single progenitor radial velocity can constrain the dark matter halo mass ($M_{200}>4.7\times 10^{12}~M_{\odot}$). Additionally, radial velocity and distance gradients along the stream offer further mass constraints. However, a more exhaustive parameter search over both progenitor and halo properties is still needed, and was not performed in Pearson et al. (2022b) due to computational limitations of their approach.
@@
In spherical dark matter halo potentials, where the orbital plane itself does not precess, the precession per orbit between apocenter or pericenter is determined by both the orbital properties (e.g., ratio of radial to azimuthal periods; Johnston et al. 2001; Binney and Tremaine 2008; Hendel and Johnston 2015), and the slope of the potential (Belokurov et al., 2014). Walder et al. (2025) recently showed that for an orbit launched at apocenter and viewed at a 0 degree inclination, multiple wraps of the orbit can constrain the radial profile of the dark matter halo.
@@
The paper is organized as follows: In §II, we provide an overview of stellar stream evolution in potentials with different radial profiles. In §III, we introduce our method and lay out the details of the sampling technique. In §IV, we show the results of applying our method to distinct streams on different orbits, in §V we discuss our findings, and in §VI we summarize and conclude.
II Stream morphologies as a probe of radial profiles
@@
where $M_{\rm halo}$ is the halo scale mass, $r_{s}$ is the scale radius of the potential, $c_{\rm NFW}$ is the halo concentration, $\gamma$ is the slope of the inner density profile, and $\beta$ is the slope of the density profile beyond the scale radius. The exponential cutoff ensures a finite mass. We use $r_{\rm cut}=25\times r_{s}$, and solve Poisson’s equation on a radial grid to infer an interpolated potential from the density. This choice of $r_{\rm cut}$ is not crucial to our modeling, since all streams have apocenters much lower than the minimum $r_{\rm cut}$ used (i.e., for $r_{s,\rm{min}}=10~\rm{kpc}$, $r_{\rm cut}=250~\rm{kpc}$). In Fig. 1 we show examples of three different dark matter halo density profiles: (1) An NFW profile (Navarro et al., 1997) with $(\gamma$$\beta)$ = ($1,3$; solid line), (2) a cored profile, with $(\gamma$$\beta)$ = ($0.2,3$; dotted line), and (3) an inner slope following an NFW profile, but a shallower outer slope $(\gamma$$\beta)$ = ($1,2$; dashed line).
@@
Table 1: Parameters for the inner and outer streamfootnotetext: Note that $r_{\rm peri}$ and $r_{\rm apo}$ are listed for the six different streams shown in Fig. 2 evolved in the three different potentials shown in Fig. 1. The first values for each stream are for the fiducial NFW halo.
@@
To illustrate the sensitivity of streams to variations in halo density profiles, in Fig. 2 we show two different streams which we label the inner stream (top row) and the outer stream (bottom row). The streams are evolved in the three different potentials shown in Fig. 1 (left: NFW, middle: cored, right: shallower outer profile), with a halo scale mass defined in Eq. 2 of $M_{\rm halo}=10^{13}$ M (corresponding to $M_{\rm enc}$ at the outer stream’s radius of $\approx 4\times 10^{12}$), $r_{\rm s,halo}=22$ kpc, $c_{\rm NFW}=15$, including a disk (Miyamoto and Nagai, 1975) with $m_{\rm disk}=5\times 10^{10}$ M, $a_{\rm disk}=3$ kpc, $b_{\rm disk}=0.2$ kpc. Our halo mass is higher than typical spiral galaxies in the local group, but compatible with lower limits from Centaurus A (Pearson et al., 2022b)  [^1] Note that the density structure of faint dwarf galaxies ($M_{*}\sim 10^{6}$ M) would be of particular interest, as stellar feedback in such galaxies would not modify the central structure (Bullock and Boylan-Kolchin, 2017).. We list the initial conditions and properties of each of the streams in Table 1. Note that throughout the paper, we fix the disk parameters, but we vary the halo parameters (see Table 2). The red stars in the upper panels show the progenitor location, and the red dashed line shows the past 500 Myr of evolution for each progenitor. Note that the streams were evolved for 5 Gyr.
@@
[…]
@@
The top two panels of Fig. 3 show the angle between successive apocenters (i.e., the angle between orbital turning points), $\theta_{\rm apo}$, for the streams evolved under the mass model in Eq. 1 with varying $\gamma$ and $\beta$ (see labels) as a function of galactocentric radius. We calculate $\theta_{\rm apo}$ using the epicycle approximation (Appendix A). In Fig. 3, steeper inner and outer density slopes (larger $\gamma$ and larger $\beta$) lead to smaller angles between successive apocenters. Note, however, that at small galactocentric radii $\theta_{\rm apo}$ is more sensitive to changes in $\gamma$, whereas at larger galactocentric radii $\theta_{\rm apo}$ is more sensitive to changes in $\beta$ due to the relations between the radial and azimuthal periods (see Appendix A). We can understand the trends between $(\gamma,\beta)$ and $\theta_{\rm apo}$ as follows. As $\gamma$ increases, the inner regions of the potential become deeper and the radial and azimuthal periods both decrease, with their ratio approaching unity. Thus, the angle between apocenters diminishes. For $\beta$, the mass enclosed approximately goes like $M_{\rm enc}\propto r^{3-\beta}$ below the cutoff radius, $r<r_{\rm cut}$, though beyond the scale radius, $r_{s}$. At large $r$, increasing to $\beta>3$ the mass distribution becomes more centrally packed-in and orbits become approximately Keplerian, for which $\theta_{\rm apo}\xrightarrow[]{}0$. Thus, we expect that increasing $\beta$ causes turning point angles to decrease. Indeed, this is what we find in Fig. 2, and show analytically in Fig. 3. We note that the angles between turning points are degenerate with projection effects. The generative method we develop in §III will account for viewing angle degeneracies.
@@
We have demonstrated that two different extragalactic streams at two different locations in their host halo are sensitive to different parts of the potential (as expected from Milky Way studies, e.g., Bonaca and Hogg, 2018; Nibauer et al., 2022), and that streams, which cross or are within the scale radius, are particularly sensitive to the inner slope of the potential. Note that while deviations from spherical symmetry can also introduce orbital plane precession, in this paper we focus on spherical dark matter halos as the simplest mass model that is also interesting for dark matter physics (Spergel and Steinhardt, 2000; Bullock and Boylan-Kolchin, 2017).
@@ III Method
III.1 Stream Generation and Model Parameters
To generate stellar streams in different potential models, we use the particle-spray method (Fardal et al., 2015) implemented in the GPU accelerated code streamsculptor (Nibauer et al., 2025), built within the JAX python framework (Bradbury et al., 2018). The GPU support of our simulator allows us to efficiently generate stream models in batches of $500$, where each stream in the batch is simulated within a different underlying potential. Each model stream consists of 1000 particles, which is sufficient for the low resolution data representation that we will discuss in §III.2. Despite the lack of dynamical friction and star-by-star force calculations, the particle spray approach successfully characterizes the tracks of dwarf galaxy streams when compared to NN-body simulations (Fardal et al., 2015; Pearson et al., 2022b).
We consider the two streams presented in §II, which trace different locations of the same potential. We aim to investigate whether the full radial profile of the host halo (i.e., the inner and outer density slopes) can be recovered from the morphology of the two streams. In Table 2 we list all the free and fixed parameters adopted in our sampling.
To generate streams, we fix the on-sky location of the progenitors. Several extragalactic streams have tentative progenitor locations (see e.g., Pearson et al. 2022b; Miró-Carretero et al. 2024a). However, future work can add free parameters for the progenitor’s on-sky location (e.g., along the ridgeline of the stream’s track). We find that incorrectly choosing the angular location of the progenitor biases progenitor parameters, but does not change our conclusions on inferring the full radial density profile (inner and outer) of the halo. We fit for the line-of-slight location of the progenitor in its host galaxy, $y_{\rm prog}$, along with its present-day velocity direction, sampled on the velocity unit sphere (see §II and Table 2). When sampling from our prior, we test whether the progenitor’s orbit is bound. We reject parameters that produce unbound orbits.